### v2

@@ Abstract
Keywords:
Vehicle trajectory prediction Multi-dataset framework Domain generalization
@@ 1 Introduction
Exploring these two research questions involves leveraging multiple trajectory prediction datasets. Firstly, these datasets provide diverse domains, allowing for a thorough examination of model generalization across different domains (RQ1). Secondly, combining these datasets creates a much larger dataset, enabling an exploration of the asymptotic limits of data scaling (RQ2). However, significant challenges exist when attempting to leverage multiple datasets. (1) Each of these datasets has a unique data format, posing practical difficulties for researchers utilizing multiple datasets. (2) Each of the datasets undergoes collection and annotation through distinct strategies, with semi-automatic pre-annotations and manual curations [7, 14, 8]. This leads to multiple discrepancies such as variations in resolution, sampling rates, and types of semantic annotations. (3) Comparing model performance across datasets is not straightforward due to varying dataset settings (e.g., prediction horizons) and evaluation metrics (e.g., mAP metric is used in WOMD [14] and brier-FDE metric in Argoverse 2 [50]). In short, while each of the datasets contributes to the progress in the field, they have been developed independently, without considering harmonization with existing ones. As a result, many trajectory prediction studies train and evaluate their models using a single dataset [13, 31, 2, 9, 46, 3, 37, 20, 22, 5].
@@
To tackle these challenges, we introduce ‘UniTraj’, a comprehensive vehicle trajectory prediction framework. UniTraj seamlessly integrates and unifies multiple data sources (including nuScenes [7], Argoverse 2 [50] and Waymo Open Motion Dataset - WOMD  [14]), models (including AutoBot [17], MTR [46], and Wayformer [35]), and evaluations. UniTraj not only serves as a solution to tackle our research questions but also provides a comprehensive and flexible platform for the community. First, it is designed for the effortless inclusion of new datasets by proposing a unified data structure compatible with various datasets. Second, Unitraj supports and simplifies the integration of new methods by providing numerous essential data processing and loss functions relevant to the trajectory prediction task. Lastly, UniTraj offers unified evaluation metrics, as well as diverse and insightful evaluation approaches, such as analyzing performance on the long-tail data instances as well as different clusters of data samples to allow a more in-depth understanding of model behavior. Figure 1 shows the overview of the framework.
@@
We conduct extensive experiments using the UniTraj framework to shed light on our two research questions. Our findings reveal a large performance drop when transitioning between data sources, alongside variations in the generalization abilities induced by different datasets (RQ1). We also show that scaling up dataset size and diversity can enhance model performance significantly without any architectural modifications, leading us to rank $1$st in the nuScenes public leaderboard. This is accomplished by training models on all existing datasets in the framework. This unified dataset forms the largest public data one can use to train a vehicle trajectory prediction model, with more than $2$M samples, $1337$ hours of data, and $15$ different cities. Finally, by providing an in-depth analysis of the datasets, we offer a more comprehensive understanding of their characteristics. Our analysis reveals that the datasets’ generalization capabilities are not only attributed to their size, but also their intrinsic diversity. We believe that the framework opens up new opportunities in the trajectory prediction field, and we will release the framework to foster further advancements. In summary, our contributions are as follows:
@@ 2 Previous work
Trajectory prediction benchmarks. Multi-dataset benchmarks have already been explored in various domains such as object detection [57], semantic segmentation [24, 45] and action recognition [30]. In the field of trajectory prediction, such benchmarks have primarily been developed for human trajectory prediction [40, 1, 39]. Notably, Trajnet++ [26] provides an interaction-centric benchmark by categorizing trajectories based on the presence of an interaction. trajdata [21] is a unified interface to multiple human trajectory datasets incorporating scene context into the inputs. A related work for the vehicle trajectory planning task is ScenarioNet [28], a simulator aggregating multiple real-world datasets into a unified format and providing a planning development and evaluation framework. We leverage their data conversion to process datasets. To the best of our knowledge, we are the first to propose an open-source framework for vehicle trajectory prediction. Our framework is not limited to including multiple datasets; it also integrates a variety of trajectory prediction models and evaluation methodologies, thereby providing a comprehensive resource for advancing research and development in the vehicle trajectory prediction task.
Generalization of trajectory prediction models. The discrepancies in data formats in vehicle trajectory prediction datasets hinder research on cross-dataset generalization, leading to limited studies in this area. In [53], one dataset is divided into different domains to explore model generalization. Authors in [44] propose an epistemic uncertainty estimation approach and perform cross-dataset evaluation. In [16], the authors studied cross-dataset generalization of models and showed a performance gap between datasets. However, they provide limited insights into the sources of the generalization gap. Moreover, their code is not publicly available. Previous works also investigated some generalization aspects of trajectory prediction models when they deal with new scenes and cities [11, 27, 32], new agent types [27], using perception outputs instead of curated annotations [49, 56, 51], and facing adversarial situations [3, 41]. In this work, we conduct more extensive and in-depth cross-dataset, and cross-city analyses as well as multi-dataset training. Moreover, we provide insights into the dataset characteristics, explaining the findings. We also release an open-source framework to facilitate this line of research.
@@ 3 UniTraj framework
The UniTraj framework, illustrated in Figure 1, consists of three main components. The first component unifies the format and features of various datasets (see Section 3.1). The second component adapts trajectory prediction models to the unified data format, facilitating their training (see Section 3.2). The final component consists of a comprehensive and shared evaluation process for the models (see Section 3.3). The integration of the components allows for diverse experimentation, such as cross-dataset training, evaluation, and dataset analysis.
@@ 3.1.2 Unified data features:
Despite the data being converted into a unified format, significant discrepancies persist across the datasets, affecting model performance. For example, the scenarios are $11$ seconds long in Argoverse 2, while they are $9$ seconds in WOMD; or the precision of map annotations are $1$ meters in nuScenes while they are $0.5$ meters in WOMD. Therefore, we aim to harmonize these discrepancies and minimize their impact on the model’s performance. Table 1 summarizes the discrepancies and the unified features. Our data processing approach involves specific harmonizations, including the following:
@@
  • Time Length. The trajectories in different datasets are with the same frequency of $10$Hz but with a duration ranging from 8 to 20 seconds. To standardize this aspect, we truncate all trajectories to a uniform length of 8 seconds. Within this duration, UniTraj provides the option to flexibly determine a unified length of past and future trajectories for all datasets.
@@ 3.2 Unified models
The models’ capacities cover a large range ($1.5$M parameters for AutoBot, $16.5$M for MTR, and $60.1$M for Wayformer), enabling research on model size and scaling.
@@ 4 Experiments
The UniTraj framework opens up new opportunities for research and experimentation. This section presents experiments highlighting these opportunities, focusing on cross-domain (i.e., cross-dataset and cross-city) generalization (RQ1) in Section 4.1, and data scaling impact (RQ2) for trajectory prediction models in Section 4.2. We provide fine-grained dataset analyses and discussions in Section 4.3. Additional experiments in the appendix, such as continual learning and synthetic-to-real transfer, further demonstrate the framework’s research utility.
Experimental settings: We replicate the model configurations and hyper-parameters from their original implementations. Throughout the experiments, we have limited the training and validation samples to vehicle trajectories. The map range extends to a $100$m radius with a spatial resolution of $0.5$m. The temporal parameters are set to $2$ seconds of historical trajectories and $6$ second future trajectories. For our multi-dataset training experiments, we utilize WOMD [14], Argoverse 2 [50], and nuScenes [7] datasets. Since the nuPlan [8] dataset is oriented towards planning tasks and lacks an official training/validation set for prediction tasks, we exclusively use it for the cross-city generalization studies due to its large number of samples for different cities. We only report the results with the brier-minFDE metric and leave other metrics for the appendix.
@@ 4.1.1 Cross-dataset evaluation:
To assess the generalization capabilities of models, we train models on each individual dataset and evaluate their performance on all other available datasets. The findings are presented in Table 2. Analyzing the data in different columns of the table, the first observation is that all models’ performances decline significantly when models are tested on other datasets. This is a consistent trend across all of the three model architectures, and all of the considered datasets. For instance, the second column under MTR reports the performance evaluated on the validation set of Argoverse 2. It indicates that MTR achieves its peak performance when it is trained on the training set of Argoverse 2 itself, while models trained on nuScenes and WOMD exhibit significantly lower performances.
@@
nuScenes 32k 3.06 4.67 7.53 3.85 4.93 7.74 3.36 4.48 6.89
Argoverse 2 180k 3.89 2.31 4.84 5.21 2.85 5.43 4.35 2.51 4.43
WOMD 1800k 3.28 3.84 2.37 3.75 3.64 2.64 3.73 3.23 2.47
All 2012k 2.49 2.22 2.37 3.34 2.56 2.61 3.07 2.54 2.47
@@ 4.1.2 Cross-city evaluation:
Despite our care to standardize data formats and align features among datasets, certain fundamental discrepancies may persist caused by the data collection and annotation processes. For instance, annotation noises could still exist across datasets. To control for this potential residual discrepancy, we explore the generalization of AutoBot when the city is changed inside a single dataset. Similar to the previous experiment, we train AutoBot on each city and evaluate it on the rest of cities. We employed nuPlan [8] data for this experiment due to the large number of samples existing in diverse cities and selected $10$K samples from each city. The results are shown in Table 3. It shows that the performance of AutoBot drops once evaluated on other cities. For instance, the first row shows that the model trained on Pittsburgh has the best performance on Pittsburgh (brier-minFDE $2.4$) and worse performances on Boston ($2.7$) and Singapore ($3.5$). This indicates a clear generalization gap between cities, emphasizing the discrepancies between different environments. Moreover, it can also be observed that the model trained on Singapore performs the worst on average.
@@
The unified data available in UniTraj forms the largest public data one can use to train a trajectory prediction model. In this section, we explore if we can improve the models’ performance by simply scaling the size of the training dataset. Therefore, we combine all the existing real datasets in UniTraj into a single large training set on which we train the considered models. The results of this experiment are presented at the bottom of Table 2 (row ‘All’), demonstrating improvements over the model trained solely on a single dataset. While the improvements are not identical for different datasets, they are particularly significant in the case of the nuScenes dataset, making the MTR model trained on combined data rank $1$st in the nuScenes leaderboard (shown in Section 4.2). For instance, training the MTR model on "All" datasets enables it to outperform the model trained on nuScenes and Argoverse2 by a large margin. Moreover, while the performance on WOMD has not been improved, the resulting model performs much better than the model trained on WOMD on other datasets. These improvements are attributed to the relatively larger size and diversity of the combined dataset compared to each individual one. We elaborate on this more in Section 4.3.
The table also shows that certain models benefit more from larger data sizes compared to others. Specifically, when looking at the performance on Argoverse 2 and nuScenes datasets, which benefit from the increased dataset size, models like MTR and Wayformer show more significant improvements than AutoBot. This difference in performance enhancement is attributed to the models’ capacity. For instance, MTR has $60$ M parameters, providing it with a higher capacity to learn from larger datasets, whereas AutoBot, with only $1$ million parameters, may not be as able to utilize the additional data.
In order to illustrate the impact of data size on the performance of a trajectory predcition model, we gradually increase the number of training samples from $20$% to $100$% of the combined dataset. We then report the AutoBot model’s performance using the average brier-minFDE metric among all three datasets. Section 4.2 shows the curve revealing a consistent reduction in the prediction error once the dataset size increases. This highlights the substantial benefits of larger datasets on the model’s performance and offers prospects for improved performances with larger data sizes.
@@
Method Ranking ($\downarrow$) minADE5555 ($\downarrow$)
MTR-UniTraj 1 0.98
@@ 4.3 Analyzing the results
In this section, we aim to explain the findings related to generalization gaps and the data scaling impact observed in Sections 4.1.1 and 4.2. To do so, we first delve into a comparative analysis of datasets integrated within the UniTraj framework with our fine-grained evaluations. We then employ these insights to understand the findings.
@@
StationaryStraightStraight RightStraight LeftRight TurnLeft TurnLeft U-TurnRight U-Turn00202020204040404060606060# trajectories (%)WOMDArgoverse 2nuScenes
@@
0–1010–2020–3030–4040–5050–6060–7070–80002020202040404040Refer to captionRefer to captionRefer to captionKalman difficulty# trajectories (%)WOMDArgo. 2nuScenes
@@
Trajectory type based comparison: The analysis of trajectory types in the WOMD, Argoverse 2, and nuScenes datasets in Figure 3(a) reveals trajectory type imbalances, primarily featuring a prevalence of straight trajectories, constituting 54% to 68% of all trajectories, with minimal instances of u-turns. WOMD exhibits a notably diverse trajectory mix, with a significant number of left and right turn trajectories, approximately two times more than what is observed in Argoverse 2 and nuScenes. This stands in contrast to Argoverse 2 and nuScenes, which primarily contain straight trajectories over varied turning maneuvers.
Kalman difficulty based comparison: The distribution of sample difficulties within the WOMD, Argoverse 2, and nuScenes datasets, shown in Figure 3(b), exhibits a consistent trend where easier scenarios significantly outnumber more challenging ones. WOMD demonstrates a relatively balanced distribution across lower to moderate difficulty levels, with around $\sim$24of trajectories falling within the easiest category (Kalman difficulties up to 10), with a consistent presence observed up to a Kalman difficulty of 50.0. In contrast, both nuScenes and Argoverse 2 exhibit a substantial bias towards easier difficulties, comprising approximately $\sim$42and $\sim$46of samples, respectively, in the lowest difficulty range (Kalman difficulties lower than 10), and show a sharp decrease in proportion with increasing difficulty levels, indicating datasets primarily composed of simpler scenarios which may potentially limit their efficacy in training models for more complex situations.
@@
MTR (nuScenes) 2.15 2.58 4.85 4.26 8.13 4.82 5.17 4.85 3.06
MTR (All) 2.23 2.31 3.13 3.06 2.98 3.53 2.10 2.82 2.49
Table 6: Cross-dataset generalization experiments with identical sample size. We select $30$K random samples from every dataset, then train and validate AutoBot across them and report the brier-minFDE metric.
@@
$\downarrow$ Training Argoverse 2 WOMD nuScenes Average
Argoverse 2 2.90 5.07 4.67 4.21
WOMD 4.04 3.22 4.42 3.89
nuScenes 4.48 6.88 3.38 4.91
All 2.80 3.13 3.25 3.06
@@
difficulty [0,30[$\in$[0,30[∈ [ 0 , 30 [ [30,60[$\in$[30,60[∈ [ 30 , 60 [ [60,100[$\in$[60,100[∈ [ 60 , 100 [
MTR (nuScenes) 2.93 4.72 5.01
MTR (All) 2.43 3.17 7.86
Explaining the findings in cross-dataset generalization and multi-dataset training experiments: Our cross-dataset generalization experiment in Section 4.1.1 shows that models do not generalize equally across different datasets. Notably, models trained on WOMD generalize better to other datasets. Furthermore, models trained on combined datasets exhibit considerable improvements. To understand these phenomena, it’s important to delve into the differences between datasets, focusing on two main aspects: size and diversity.
To investigate the impact of dataset size, we replicate the cross-dataset generalization experiments (Table 2), but with control on the dataset size, as we select $30$k random samples for each dataset’s training set. Table 7 shows the results. The last column shows the generalization hierarchy where again the WOMD generalizes best, followed by Argoverse 2 and then nuScenes. This shows that better cross-dataset generalization is not solely attributed to the size of the datasets. The last row illustrates that for multi-dataset training, there is a considerable improvement for all the datasets, highlighting the pronounced benefit of adding more data in small-scale dataset scenarios.
Section 4.3.1 reveals that the datasets are dissimilar in terms of diversity. Notably, WOMD encompasses the most diverse range of scenarios in comparison to other datasets. This explains the superior generalization of WOMD to other datasets, as the diversity enables models to learn the full spectrum of data distributions more comprehensively. Similarly, the combined dataset provides a more diverse collection of trajectories, leading to enhanced performance for the models. To demonstrate this, we compare the fine-grained evaluations of MTR model trained on nuScenes and the combined dataset. Table 5 shows the per trajectory type performance of the two models where the model trained on full data outperforms in every trajectory type since the full data includes significantly more samples from each trajectory type. We also compare the performances using the Kalman difficulty measure in Table 7. The combined data has considerably more medium-difficulty samples (shown in Figure 3(b)) leading to significant performance improvements in the medium-range samples. We also note a performance drop in hard examples due to the insufficient challenging scenarios in the datasets. These results highlight the importance of diversity in the data.

### v3

@@ 1 Introduction
Exploring these two research questions involves leveraging multiple trajectory prediction datasets. Firstly, these datasets provide diverse domains, allowing for a thorough examination of model generalization across different domains (RQ1). Secondly, combining these datasets creates a much larger dataset, enabling an exploration of the asymptotic limits of data scaling (RQ2). However, significant challenges exist when attempting to leverage multiple datasets. (1) Each of these datasets has a unique data format, posing practical difficulties for researchers utilizing multiple datasets. (2) Each of the datasets undergoes collection and annotation through distinct strategies, with semi-automatic pre-annotations and manual curations [7, 14, 8]. This leads to multiple discrepancies such as variations in resolution, sampling rates, and types of semantic annotations. (3) Comparing model performance across datasets is not straightforward due to varying dataset settings (e.g., prediction horizons) and evaluation metrics (e.g., mAP metric is used in WOMD [14] and brier-FDE metric in Argoverse 2 [52]). In short, while each of the datasets contributes to the progress in the field, they have been developed independently, without considering harmonization with existing ones. As a result, many trajectory prediction studies train and evaluate their models using a single dataset [13, 30, 2, 9, 48, 3, 37, 20, 22, 5, 34].
@@
To tackle these challenges, we introduce ‘UniTraj’, a comprehensive vehicle trajectory prediction framework. UniTraj seamlessly integrates and unifies multiple data sources (including nuScenes [7], Argoverse 2 [52] and Waymo Open Motion Dataset - WOMD  [14]), models (including AutoBot [17], MTR [48], and Wayformer [35]), and evaluations. UniTraj not only serves as a solution to tackle our research questions but also provides a comprehensive and flexible platform for the community. First, it is designed for the effortless inclusion of new datasets by proposing a unified data structure compatible with various datasets. Second, Unitraj supports and simplifies the integration of new methods by providing numerous essential data processing and loss functions relevant to the trajectory prediction task. Lastly, UniTraj offers unified evaluation metrics, as well as diverse and insightful evaluation approaches, such as analyzing performance on the long-tail data instances as well as different clusters of data samples to allow a more in-depth understanding of model behavior. Figure˜1 shows the overview of the framework.
@@
We conduct extensive experiments using the UniTraj framework to shed light on our two research questions. Our findings reveal a large performance drop when transitioning between data sources, alongside variations in the generalization abilities induced by different datasets (RQ1). We also show that scaling up dataset size and diversity can enhance model performance significantly without any architectural modifications, leading us to rank 11st in the nuScenes public leaderboard. This is accomplished by training models on all existing datasets in the framework. This unified dataset forms the largest public data one can use to train a vehicle trajectory prediction model, with more than 22M samples, $1337$ hours of data, and $15$ different cities. Finally, by providing an in-depth analysis of the datasets, we offer a more comprehensive understanding of their characteristics. Our analysis reveals that the datasets’ generalization capabilities are not only attributed to their size, but also their intrinsic diversity. We believe that the framework opens up new opportunities in the trajectory prediction field, and we will release the framework to foster further advancements. In summary, our contributions are as follows:
@@ 2 Previous work
Trajectory prediction benchmarks. Multi-dataset benchmarks have already been explored in various domains such as object detection [59], semantic segmentation [24, 47], and pose prediction [41]. In the field of trajectory prediction, such benchmarks have primarily been developed for human trajectory prediction [42, 1, 39]. Notably, Trajnet++ [26] provides an interaction-centric benchmark by categorizing trajectories based on the presence of an interaction. trajdata [21] is a unified interface to multiple human trajectory datasets incorporating scene context into the inputs. A related work for the task of vehicle trajectory planning is ScenarioNet [28], a simulator aggregating multiple real-world datasets into a unified format and providing a planning development and evaluation framework. To the best of our knowledge, we are the first to propose an open-source framework for vehicle trajectory prediction. Our framework is not limited to including multiple datasets; it also integrates a variety of trajectory prediction models and evaluation methodologies, thereby providing a comprehensive resource for advancing research and development in the vehicle trajectory prediction task.
Generalization of trajectory prediction models. The discrepancies in data formats in vehicle trajectory prediction datasets hinder research on cross-dataset generalization, leading to limited studies in this area. In [55], one dataset is divided into different domains to explore model generalization. Authors in [46] propose an epistemic uncertainty estimation approach and perform cross-dataset evaluation. In [16], the authors studied cross-dataset generalization of models and showed a performance gap between datasets. However, they provide limited insights into the sources of the generalization gap. Moreover, their code is not publicly available. Previous works also investigated some generalization aspects of trajectory prediction models when they deal with new scenes and cities [11, 27, 31], new agent types [27], using perception outputs instead of curated annotations [51, 58, 53], and facing adversarial situations [3, 43, 40]. In this work, we conduct more extensive and in-depth cross-dataset, and cross-city analyses as well as multi-dataset training. Moreover, we provide insights into the dataset characteristics, explaining the findings. We also release an open-source framework to facilitate this line of research.
@@ 3 UniTraj framework
The UniTraj framework, illustrated in Figure˜1, consists of three main components. The first component unifies the format and features of various datasets (see Section˜3.1). The second component adapts trajectory prediction models to the unified data format, facilitating their training (see Section˜3.2). The final component consists of a comprehensive and shared evaluation process for the models (see Section˜3.3). The integration of the components allows for diverse experimentation, such as cross-dataset training, evaluation, and dataset analysis.
@@ 3.1.2 Unified data features:
Despite the data being converted into a unified format, significant discrepancies persist across the datasets, affecting model performance. For example, the scenarios are $11$ seconds long in Argoverse 2, while they are $9$ seconds in WOMD; or the precision of map annotations are $1$ meters in nuScenes while they are $0.5$ meters in WOMD. Therefore, we aim to harmonize these discrepancies and minimize their impact on the model’s performance. Table˜1 summarizes the discrepancies and the unified features. Our data processing approach involves specific harmonizations, including the following:
@@
  • Time Length. The trajectories in different datasets are with the same frequency of 1010Hz but with a duration ranging from 8 to 20 seconds. To standardize this aspect, we truncate all trajectories to a uniform length of 8 seconds. Within this duration, UniTraj provides the option to flexibly determine a unified length of past and future trajectories for all datasets.
@@ 3.2 Unified models
The models’ capacities cover a large range (1.51.5M parameters for AutoBot, 60.160.1M for MTR, and 16.516.5M for Wayformer), enabling research on model size and scaling.
@@ 4 Experiments
The UniTraj framework opens up new opportunities for research and experimentation. This section presents experiments highlighting these opportunities, focusing on cross-domain (i.e., cross-dataset and cross-city) generalization (RQ1) in Section˜4.1, and data scaling impact (RQ2) for trajectory prediction models in Section˜4.2. We provide fine-grained dataset analyses and discussions in Section˜4.3. Additional experiments in the appendix, such as continual learning and synthetic-to-real transfer, further demonstrate the framework’s research utility.
Experimental settings: We replicate the model configurations and hyper-parameters from their original implementations. Throughout the experiments, we have limited the training and validation samples to vehicle trajectories. The map range extends to a 100100m radius with a spatial resolution of 0.50.5m. The temporal parameters are set to $2$ seconds of historical trajectories and $6$ second future trajectories. For our multi-dataset training experiments, we utilize WOMD [14], Argoverse 2 [52], and nuScenes [7] datasets. Since the nuPlan [8] dataset is oriented towards planning tasks and lacks an official training/validation set for prediction tasks, we exclusively use it for the cross-city generalization studies due to its large number of samples for different cities. We only report the results with the brier-minFDE metric and leave other metrics for the appendix.
@@ 4.1.1 Cross-dataset evaluation:
To assess the generalization capabilities of models, we train models on each individual dataset and evaluate their performance on all other available datasets. The findings are presented in Table˜2. Analyzing the data in different columns of the table, the first observation is that all models’ performances decline significantly when models are tested on other datasets. This is a consistent trend across all of the three model architectures, and all of the considered datasets. For instance, the second column under MTR reports the performance evaluated on the validation set of Argoverse 2. It indicates that MTR achieves its peak performance when it is trained on the training set of Argoverse 2 itself, while models trained on nuScenes and WOMD exhibit significantly lower performances.
@@
nuScenes 32k 2.86 4.50 7.38 3.06 4.68 7.16 3.36 4.48 6.89
Argoverse 2 180k 3.72 2.08 4.68 3.69 2.38 4.80 4.35 2.51 4.43
WOMD 1800k 3.10 3.63 2.13 3.12 3.60 2.10 3.73 3.23 2.47
All 2012k 2.27 1.99 2.13 2.32 2.12 2.09 3.07 2.54 2.47
@@ 4.1.2 Cross-city evaluation:
Despite our care to standardize data formats and align features among datasets, certain fundamental discrepancies may persist caused by the data collection and annotation processes. For instance, annotation noises could still exist across datasets. To control for this potential residual discrepancy, we explore the generalization of AutoBot when the city is changed inside a single dataset. Similar to the previous experiment, we train AutoBot on each city and evaluate it on the rest of cities. We employed nuPlan [8] data for this experiment due to the large number of samples existing in diverse cities and selected 1010K samples from each city. The results are shown in Table˜3. It shows that the performance of AutoBot drops once evaluated on other cities. For instance, the first row shows that the model trained on Pittsburgh has the best performance on Pittsburgh (brier-minFDE $2.4$) and worse performances on Boston ($2.7$) and Singapore ($3.5$). This indicates a clear generalization gap between cities, emphasizing the discrepancies between different environments. Moreover, it can also be observed that the model trained on Singapore performs the worst on average. This is an expected outcome, given that Singapore is a left-hand traffic city, unlike the other ones.
@@
The unified data available in UniTraj forms the largest public data one can use to train a trajectory prediction model. In this section, we explore if we can improve the models’ performance by simply scaling the size of the training dataset. Therefore, we combine all the existing real datasets in UniTraj into a single large training set on which we train the considered models. The results of this experiment are presented at the bottom of Table˜2 (row ‘All’), demonstrating improvements over the model trained solely on a single dataset. While the improvements are not identical for different datasets, they are particularly significant in the case of the nuScenes dataset, making the MTR model trained on combined data rank 11st in the nuScenes leaderboard (shown in Section˜4.2). For instance, training the MTR model on "All" datasets enables it to outperform the model trained on nuScenes and Argoverse2 by a large margin. Moreover, while the performance on WOMD has not been improved, the resulting model performs much better than the model trained on WOMD on other datasets. These improvements are attributed to the relatively larger size and diversity of the combined dataset compared to each individual one. We elaborate on this more in Section˜4.3.
The table also shows that certain models benefit more from larger data sizes compared to others. Specifically, when looking at the performance on Argoverse 2 and nuScenes datasets, which benefit from the increased dataset size, models like MTR and Wayformer show more significant improvements than AutoBot. This difference in performance enhancement is attributed to the models’ capacity. For instance, MTR has $60.1$ M parameters, providing it with a higher capacity to learn from larger datasets, whereas AutoBot, with only $1.5$ M parameters, may not be as able to utilize the additional data.
In order to illustrate the impact of data size on the performance of a trajectory predcition model, we gradually increase the number of training samples from 2020% to 100100% of the combined dataset. We then report the AutoBot model’s performance using the average brier-minFDE metric among all three datasets. Section˜4.2 shows the curve revealing a consistent reduction in the prediction error once the dataset size increases. This highlights the substantial benefits of larger datasets on the model’s performance and offers prospects for improved performances with larger data sizes.
@@
Method Ranking ($\downarrow$) minADE55 ($\downarrow$)
MTR-UniTraj 1 0.96
@@ 4.3 Analyzing the results
In this section, we aim to explain the findings about generalization gaps and the data scaling impact in Sections˜4.1.1 and 4.2. Thus, we first delve into a comparative analysis of datasets integrated within the UniTraj framework with our fine-grained evaluations. We then employ these insights to explain the findings.
@@
StationaryStraightStraight RightStraight LeftRight TurnLeft TurnLeft U-TurnRight U-Turn0202040406060# trajectories (%)WOMDArgoverse 2nuScenes
@@
0–1010–2020–3030–4040–5050–6060–7070–80020204040Refer to captionRefer to captionRefer to captionKalman difficulty# trajectories (%)WOMDArgo. 2nuScenes
@@
Trajectory type based comparison: The analysis of trajectory types in the WOMD, Argoverse 2, and nuScenes datasets in Figure˜3(a) reveals trajectory type imbalances, primarily featuring a prevalence of straight trajectories, constituting 54% to 68% of all trajectories, with minimal instances of u-turns. WOMD exhibits a notably diverse trajectory mix, with a significant number of left and right turn trajectories, approximately two times more than what is observed in Argoverse 2 and nuScenes. This stands in contrast to Argoverse 2 and nuScenes, which primarily contain straight trajectories over varied turning maneuvers.
Kalman difficulty based comparison: The distribution of sample difficulties within the WOMD, Argoverse 2, and nuScenes datasets, shown in Figure˜3(b), exhibits a consistent trend where easier scenarios significantly outnumber more challenging ones. WOMD demonstrates a relatively balanced distribution across lower to moderate difficulty levels, with around $\sim$24% of trajectories falling within the easiest category (Kalman difficulties up to 10), with a consistent presence observed up to a Kalman difficulty of 50.0. In contrast, both nuScenes and Argoverse 2 exhibit a substantial bias towards easier difficulties, comprising approximately $\sim$42% and $\sim$46% of samples, respectively, in the lowest difficulty range (Kalman difficulties lower than 10), and show a sharp decrease in proportion with increasing difficulty levels, indicating datasets primarily composed of simpler scenarios which may potentially limit their efficacy in training models for more complex situations.
@@
MTR (nuScenes) 2.15 2.58 4.85 4.26 8.13 4.82 5.17 4.85 2.86
MTR (All) 2.23 2.31 3.13 3.06 2.98 3.53 2.10 2.82 2.27
Table 6: Cross-dataset generalization experiments with identical sample size. We select 3030K random samples from every dataset, then train and validate AutoBot across them and report the brier-minFDE metric.
@@
$\downarrow$ Training nuScenes Argoverse 2 WOMD Average
nuScenes 3.38 4.48 6.88 4.91
Argoverse 2 4.67 2.90 5.07 4.21
WOMD 4.42 4.04 3.22 3.89
All 3.25 2.80 3.13 3.06
@@
difficulty $\in[0,30[$ $\in[30,50[$ $\in[50,100[$
MTR (nuScenes) 2.73 4.52 4.25
MTR (All) 2.23 2.97 4.20
Explaining the findings in cross-dataset generalization and multi-dataset training experiments: Our cross-dataset generalization experiment in Section˜4.1.1 shows that models do not generalize equally across different datasets. Notably, models trained on WOMD generalize better to other datasets. Furthermore, models trained on combined datasets exhibit considerable improvements. To understand these phenomena, it’s important to delve into the differences between datasets, focusing on two main aspects: size and diversity.
To investigate the impact of dataset size, we replicate the cross-dataset generalization experiments (Table˜2), but with control on the dataset size, as we select 3030k random samples for each dataset’s training set. Table˜7 shows the results. The last column shows the generalization hierarchy where again the WOMD generalizes best, followed by Argoverse 2 and then nuScenes. This shows that better cross-dataset generalization is not solely attributed to the size of the datasets. The last row illustrates that for multi-dataset training, there is a considerable improvement for all the datasets, highlighting the pronounced benefit of adding more data in small-scale dataset scenarios.
Section˜4.3.1 reveals that the datasets are dissimilar in terms of diversity. Notably, WOMD encompasses the most diverse range of scenarios in comparison to other datasets. This explains the superior generalization of WOMD to other datasets, as the diversity enables models to learn the full spectrum of data distributions more comprehensively. Similarly, the combined dataset provides a more diverse collection of trajectories, leading to enhanced performance for the models. To demonstrate this, we compare the fine-grained evaluations of MTR model trained on nuScenes and the combined dataset. Table˜5 shows the per trajectory type performance of the two models where the model trained on full data outperforms in every trajectory type since the full data includes significantly more samples from each trajectory type. We also compare the performances using the Kalman difficulty measure in Table˜7. The combined data has considerably more medium-difficulty samples (shown in Figure˜3(b)) leading to significant performance improvements in the medium-range samples. These results highlight the importance of diversity in the data.